Could you please check my answer?
Find the length of a side of a square if its area is the same as the area of a triangle with an altitude of 18 cm and a base of 11 cm. Find each answer to the nearest tenth.
My answer: 9.9 cm

Answer by Alan3354(69443)

Find the length of a side of a square if its area is the same as the area of a triangle with an altitude of 18 cm and a base of 11 cm. Find each answer to the nearest tenth.
---------------
Area of the triangle = b*h/2 = 18*11/2 = 99 sq cms
------
Area of a square = s^2 = 99 sq cm
--> s = sqrt(99)
s =~ 9.9487 cm
--> 9.9 to the nearest 10th